OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $400. How many sophomores bought tickets?

OpenStudy (anonymous):

Help <3

sam (.sam.):

F + S = 150, therefore F = 150 - S second equation is $2 x F + $3 x S = 410 or 2F + 3S = 410 We know F = 150 - S Therefore 2(150 - S) + 3S = 410 300 -2S + 3S = 410 S = 110 Substitute this back into F = 150 - S , F = 40 ------------------------------------------------- Check 40 +110 = 150 tickets 40 x $2 = $80 110 x $3 = $330 $80 + $330 = $410

OpenStudy (anonymous):

F + S = 150 tickets 2F + 3S = $400 -> F =50, S = 100 tickets

OpenStudy (anonymous):

Let s = the number of sophomores 2(150-s)+3 s=400 s=100